# SMILES has to go : Representation of Molecules via Algebraic Data Types

### v1

@@ 2.1 Types of representation
    Storage representations refer to the format of storage of raw molecular data, such as molfiles or Structured Data Files (SDF) files [Wigh et al., 2022].
@@
    Transmission representations refer to how molecules are sent between cheminformaticians and depends on the medium of communication used.
@@
    Computational representations refer to how a molecule is represented in a programming language. Examples of these are bit-vectors or matrices.
2.2 Modern string representations
As this paper is focused on modern molecular representations and their application for machine learning, the history of molecular representation and an in depth discussion of other representaions is not found here, but can be found in e.g. Krenn et al. [2022] or Wigh et al. [2022].
The types molecular representations considered here are Strings, Graphs, Fingerprints and Variational Auto Encoders, and Computation Tables, each considered in a recent review of molecular representation [Wigh et al., 2022].
2.2.1 Graphs
A molecular graph is defined as “a connected, labelled, and undirected graph. The vertices are atoms, and the edges are bonds. A vertex is labelled with an atom type (chemical element) and possibly the number of unshared electrons; an edge is labelled with a bond order or a bond type (single, double, aromatic, etc.)[Dietz, 1995]
As edges in graphs (as opposed to hypergraphs) are between two vertices only, they are limited in how they can represent delocalised and multi-centre bonding. Although it is possible to label edges with, for instance, rational bond orders, the contribution of individual electrons to multiple covalent bonds would not be described by such a graph. Attempting to express delocalised bonding by labelling edges of a molecular graph with bond types would require a sufficient number of predefined bond types, which would need to be expended for each new bonding relation not described. As Dietz said “Enhancing the expressiveness by including a new bond type for every exceptional case is certainly not a very elegant solution[Dietz, 1995]
Tautomers are unable to be expressed by molecular graphs in a single graph, instead requiring multiple structures to represent the different forms of the molecule.
Furthermore graphs detrimentally interact with tautomers, according to [David et al., 2020] as evidenced by the following: “[For] applications where one is limited to using a single static representation for a molecule that is in fact rearranging on the timescale of the problem (e.g. tautomers), then a single molecular graph representation would not be appropriate and could even be detrimental to solving the problem.”.
2.2.2 Hypergraphs
A hypergraph is a graph where edges may connect any number of vertices.
In a molecular hypergraph, edges with two verticies represent localised bonding, and edges with more than two verticies represent delocalised bonding, and are labelled with the number of electrons shared between the corresponding atoms.
From Dietz:“A hyperedge containing more than two atoms gives us no information about the binary neighborhood relationships between them. That means we have no information at our disposal conceming the way in which the electrons are delocalized over these atoms” [Dietz, 1995]. Hypergraphs would indicate an electronic “soup” with out knowledge of the manner electrons are delocalised. Consequently, they are unsuitable as as a representation.
2.2.3 Dietz representaion
The Dietz representation was designed to overcome limitations of graph based and hypergraph representation. Specifically designed as a computational representation which is as expressive as a structural formulae diagram, and which provides constitutional and stereochemistry features with ease. It has so far seen little use. It uses paddlewheels to represent 3D information, but we ignore this as not relevant so far.
In the paper, it captures diborane and ferrocene with ease, and has a general model for delocalisation, which does not require explicit specificication of bond types such as “covalent" or “aromatic".
2.2.4 SMILES
SMILES (Simplified Molecular Input Line Entry System)[Weininger, 1988] represents the structure of chemical molecules using ASCII strings. Conversion between molecular graphs and text occurs by encoding atoms and bonds with characters and symbols, following a specified grammar to indicate connectivity and structure. Despite known problems, SMILES remains the predominant contemporary representation of molecules[Gómez-Bombarelli et al., 2018], [Semenova et al., 2020] & [Deng et al., 2022]
Not all valid smiles strings correspond to valid molecules. The problem of “syntactic invalidity” means that programmers may parse technically valid smiles strings but fail to produce molecules, and machine learning algorithms can output syntactic nonsense. Machine learning algorithms must re-sample when they encounter syntactic nonsense during the input or output [Gómez-Bombarelli et al., 2018].
In SMILES, small changes in the string can lead to large changes in molecular structure and adjacent atoms in the molecule can be far apart in the string [Fang et al., 2022]. As well as impacting human understanding, this can make it difficult to write or even find a similarity metric of SMILES strings whose magnitude corresponds to a relevant similarity of molecules.
As SMILES strings correspond to molecular graphs, they inherit the weaknesses of graphs as a form of representation. SMILES strings cannot represent delocalised bonds, resonance forms, or tautomerism. Where resonance or tautomerism is represnted, separate structures representing different forms or kukule structures must represented, or chosen between. Ionic bonds, polycentric bonds and metal-metal bonds similarly cannot be represented.
SMILES strings are unable to represent multiple types of stereochemistry. Atropisomers, complex molecules with multiple stereocentres, and non-tetrahedral stereocentres cannot be represented. Without 3D information, it is not possible to represent many forms of stereochemistry, and tasks such as calculating dihedral bond angles are not possible.
Given multiple ways to represent one molecule, several problems arise. The first is that machine learning algorithms need to discern the equality of differing representations. The second is that the symmetry of a representation can be broken. Consider the symmetry of the Cyclopentadienyl anion which becomes [cH-]1cccc1 in SMILES, this does not indicate the rotational symmetry of the molecule. In particular, in SMILES no indication is given as to whether [cH-]1cccc1 = 1cccc1[cH-]. However, the Figure 2 clearly shows rotational symmetry and no signs at all of “one-sidedness”.
Figure 1: An example of a non symmetric SMILES molecule representation known as the Cyclopentadienyl anion: [cH-]1cccc1
Figure 2: Consider the complexity in reading and understanding existing well known molecules such as morphine depicted here [Dietz, 1995]
According to [Wigh et al., 2022], an issue that hasn’t even been considered with Daylight’s SMILES is that its canonicalization algorithm is proprietary, and as such implementations vary between companies and research teams [David et al., 2020]. This caused the community to standardise the implementation with OpenSMILES. Despite recent work which suggests that machine learning models which are able to output nonsensical SMILES strings perform better [Skinnider, 2024], this work has only been tested on SELFIES strings. Furthermore, as a scientific result, there is no analytic proof that these models necessarily perform better and the generality of the result “Models that generate invalid outputs outperform models that do not” is debatable, especially to cheminformaticians who are not in the machine learning community. Indeed even in the machine learning community 3D coordinate data and dihedral angles are often essential [Zhou et al., 2022], [Xu et al., 2022], [Dayalan et al., 2006] and is not included in standard SMILES strings. The title of the article is too strong for the result provided “Invalid SMILES are beneficial rather than detrimental to chemical language models”. Instead we believe the title should be specific and epistemically conservative, suggesting that SMILES strings may be better than SELFIES strings on same tasks, where SELFIES is another representation which this paper criticises. Furthermore, it has nothing to say about the readability of the molecules – and the subsequent performance increases for both bench chemists and analytic chemistry in general.
2.2.5 SELFIES
This work brings together problems with representations that previous literature reviews have failed to notice [David et al., 2020] & [Wigh et al., 2022], such as the overloading of notation or complexity of the CFG with SELFIES, or the breaking of symmetry with SMILES. Previous reviews have also left out [Srinivasan et al., 1996].
SELFIES [Krenn et al., 2020] attempts to solve problems with invalidity of SMILES strings, however, it remains unable to represent molecules with resonant structures and delocalised bonding. Furthermore, it is not easy to extend requiring significant knowledge of a complex and seemingly arbitrary grammar and as such is not easy to extend e.g. incorporating 3D atomic coordinate data. This means there is no way to calculate dihedral angles or bond lengths, which are often critical in machine learning tasks [Schütt et al., 2017]. Furthermore, it also unnaturally introduces the overloading of terms, or dual semantics for a single token, in the language. This will lead to machine learning algorithms requiring more data to discern molecular equality, not to mention difficulties in the readibility of one dimensional strings. Furthermore, there are difficulties in writing or finding a similarity metric that has guaranteed properties, guaranteed by a type and whose magnitude corresponds to a relevant similarity of molecules represented by SMILES strings. This could be problematic in using Variational Auto-encoders (VAEs) based off of SMILES strings to calculate the similarity of a molecule. This is where small changes in the representation significantly changes the properties of the molecule. Similarity metrics in the ADT proposed here can be defined transparently. The ADT also includes coordinate data and indeed, the 3D information is crucial for molecular property prediction [Zhou et al., 2022], [Xu et al., 2022].
SELFIES [Krenn et al., 2020] solves arguably the single most pressing issue with SMILES. They create a “100% ROBUST” representation, which would translate to sound and complete in computer science terminology. In particular, they present a Context Free Grammar (CFG) where all sequences of terminal symbols represent molecules (soundness) and all molecules can be represented (completeness). The CFG of SELFIES can be translated into sum and product notation as every CFG can be represented by an algebraic data type [Syrotiuk, 1984]. In addition, pattern matching, programming with or understanding their CFG would be immensely challenging, because of the significant complexity of the grammar.
Their CFG has undesirable properties from the point of view of readability and also machine learning. The authors overload standard terms in their CFG with numbers after branch or ring statements. This is seen in Figure 3. This could be related to their desire to use SELFIES with machine learning algorithms that mutate and evaluate the resulting molecule, such as genetic algorithms. However, they still wish to preserve semantic coherence under mutation, rather than having syntactic incoherence [^1] In this article, they talk about combining SELFIES with genetic algorithms: https://aspuru.substack.com/p/molecular-graph-representations-and. Figure 3 describes the overridden numeric denotation of a symbol if it appears after a branch or ring token, which breaks the straightforward readability of the grammar. The machine learning algorithm would additionally have to learn this strange pattern itself requiring more data in order to do so, and once again possibly incorrectly generalising.
Figure 3: This is how the overloading takes place. This must expand with the number of elements and types of thing in the SELFIES language. Source: [Krenn et al., 2020]
In addition, resonant structures are not directly expressible in the SELFIES language. Whilst, aromatic rings like benzene are expressible, resonant structures – which involve delocalised electrons over a set of atoms, are not generally representable. Whilst SELFIES does support rudimentary stereochemistry, SELFIES does not represent three dimensional information. It is hard to see how to easily extend the grammar ensuring that it remains to be sound. This is a problem for stereoisomers like cisplatin and transplatin, one being an approved anti-cancer drug, the other having no activity in human subjects. At this point, the soundness and completeness properties become trivial by excluding potentially important data. This could be especially relevant from the point of view of cheminformaticians, who may well need access to tautomers to solve protein-ligand binding affinity [Pospisil et al., 2003].
Furthermore, E-Z stereochemistry does not provide accurate dihedral or torsional angles [Grell et al., 2006]. Proteins rely on these angles for many tasks. In [Dayalan et al., 2006], the importance of this information for protein structure prediction is revealed: “Dihedral angles are of considerable importance in protein structure prediction as they define the backbone of a protein, which together with side chains define the entire protein conformation.”. Angles influence the energy associated with a conformation of a drug which in turn relates to the utility and stability of the drug [Anderson, 2003].
New syntaxes and grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with SMILES. Consider the widely used following syntaxes from just one vendor called Daylight222www.daylight.com. SMILES was introduced and subsequently standardised in OpenSMILES. CurlySMILES was introduced to annotate SMILES strings. CHUCKLES, its extension, CHORTLES, as well as BigSMILES were introduced to represent polymeric molecules [David et al., 2020]. SMIRKS was introduced as an extension of SMARTS to represent chemical reactions [Mobley et al., 2018] – used in the Open Reaction Database. InChi V1, not part of Daylight, was introduced and can represent basic prototautemerism [Heller et al., 2013], yet is unfortunately only probabilistically reliable. SMIRKS allows for string representation of reactions. It does not represent reaction features like temperature or pressure or catalysts. One feature it lacks is the ability to indicate the directionality of the reaction, which took a new language to be implemented ten years after the release of SMIRKS, called RinChi David et al. [2020], Wigh et al. [2022]. This was further extended years later, with another language called ProcAuxInfo, which enables the storage of metadata relating to yields, temperature, concentration, and other reaction conditions. Although, this represents definite progress, the profusion of new syntaxes and grammars, and their creative backronyms, create avoidable complexity for the cheminformatician. [Apodaca, 2022] introduces Balsa, a fully-specified language designed as a subset of SMILES. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
2.2.6 Extensions of SMILES
@@
Despite the introduction of DeepSMILES, which fixed the bracketing issue in SMILES strings, the syntactic invalidity problem was not solved completely [O’Boyle and Dalke, 2018].
2.3 Other proposed representations
2.3.1 Fingerprints
@@
Fingerprints, which typically involve encoding key features of a molecule into a vector representation, offer convenience for machine learning applications. However, a significant drawback of fingerprints is their inability to consistently represent tautomers, which are molecules that have the same atomic composition but different structural arrangements. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [Martin, 2009]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
@@
Moreover, while fingerprints can capture the presence of specific substructures within molecules, they are susceptible to bit collisions and vector sparsity, which limit their overall representational power [Fang et al., 2022]. Bit collisions occur when multiple distinct substructures are mapped to the same bit in the fingerprint, leading to ambiguity and potential false positives. Vector sparsity arises when the fingerprint representation is high-dimensional, but only a small fraction of the bits are actively used, resulting in inefficient memory usage and computational overhead.
@@
A fundamental issue with numeric vector representations, such as fingerprints, is the lack of inherent semantics associated with the individual numbers. Without additional contextual information or a predefined mapping, the numeric values in the vector do not carry intrinsic meaning. This lack of semantics can hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
@@
Furthermore, the type information associated with ADTs can guide and constrain the machine learning process. The type constraints serve as a form of prior knowledge, restricting the search space and focusing the learning algorithm on meaningful and plausible solutions. This can lead to more efficient and effective learning, as the model can leverage the structure and semantics embedded in the ADT representation.
@@
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [Capecchi et al., 2020], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint. These algebraic properties such as group theoretic properties Variational Auto Encoders (VAEs) also have problems, [Gómez-Bombarelli et al., 2018] claims obtaining structured output to be a hard and currently unsolved problem, even for a simple graph data structure.

### v2

@@ 3.8 Barriers to a general, digital, molecular representation
    The assumption of localized bonding, with standard valances
@@
    The implicit or non-explicit representation of terminal hydrogen atoms.
@@
    The inability to encode advanced stereochemical relationships.
These limitations, the authors argue (and we agree), are an inherent problem with string-based representations in their ability to represent complex bonding (such as with diborane and ferrocene) and advanced stereochemical features. Cis- and trans-platin are also given as examples of molecules current string-based representations cannot represent.
The lack of quantum chemical information, which is not currently part of any modern digital representation, is also a barrier to representation, as expresed in the Kenn paper:
Thus, it should be stressed again that in d- and f-block chemistry, as well as main-group organometallic compounds, it is often impossible to assign any particular bond orders without high-level quantum chemical calculations, due to the highly de localized nature of the bonding, where electrons are often spread out over a significant number of atoms, including the metal cen ter itself, the immediately coordinated atoms, and additional ligand groups.
Despite recognizing the need for a general digital molecular representation, capable of handling all of the above issues. The authors do not reconsider the suitability of string-based models, and instaed, the solutions presented primarily involve introducing new grammars, expanding syntax rules, and further overloading symbols, leading to increasingly complex and specialized representations.
3.9 “Molecular programming languages”
In Future Project #8, Krenn et al. (2022) propose a “molecular programming language”. They reason that strings can encode the powerful computational algorithms: “Strings can store Turing-complete programming languages: In the most general case, one can store the source code of computer programs as strings. For example, a Python file is a ‘simple string’, which is executed by the Python interpreter. Python is, of course, a Turing-complete language, which means that strings can encode the most powerful computational algorithms.”
Although the source code of a programming language can be represented in files of text, they should be distinguished from string data types. For instance, they do not support string data type operations, such as insertions, [Concatenation, Contains,indexof, Lowercase, partition, replace, reverse].
Furthermore, if the programming language is represented by a string, the question of how functions within the language are interpreted is unanswered. If they are also strings, they would at some point need to be executed by something other than a string. If they are not strings, then it remains to be understood why the authors claim that strings alone are enough to represent a Turing-Complete language, let alone the utility of such a project.
Given that Turing complete programming languages with support for recursion and logical statements already exist, it is unclear why SELFIES should be augmented to become a programming language in order to represent molecules, rather than simply using a programming language which already exists.
Future Project #9 goes beyond molecular programming languages, and proposes a programming language which is “100% robust” by finding a syntax for a programming language that however you combine elements in the instruction set, a valid program is always produced. It seems this is an attempt to generalise their own robustness of mutation to the most general data structure: programs themselves. However, it remains to be seen what sense this would have, let alone any sensible denotational semantics.
3.10 Impact on machine learning models and empirical verification
From a machine learning perspective, improving representation by expanding the grammar, creating new symbols, and overloading of terms can introduce problems, as deep learning models must implicitly learn the larger, more complicated syntax.
Unlike SMILES, VAEs trained on SELFIES have a robust latent space, where any vector can be decoded into a valid molecule. Despite this, it is not a guarantee that the latent space will be smooth, especially given the non-local encoding of graph information on strings, which both SELFIES and SMILES share.
We welcome the empirical question of the effect of overloading of tokens on machine learning models, which was raised in Future Project #2, and, in light of the result that training string models on randomised representations of molecules has on model results, suggest an experiment to determine the impact of implicit learning of the grammar in the future work section.
3.11 The Algebraic Data Type (ADT) as an Alternative
We believe that the ADT later introduced will solve all of the above mentioned problems, for representation, and for the impact on machine learning models. and will represent the general, digital molecular representation searched for, by moving beyond string-based paradigms of SMILES, SELFIES, and their numerous extensions.
3.12 Extensions of SMILES
@@
Despite the introduction of DeepSMILES, which fixed the bracketing issue in SMILES strings, the syntactic invalidity problem was not solved completely [30].
New syntaxes or grammars for string-based representation have been introduced in an attempt to represent different kinds of molecules not representable with only SMILES. For instance, there are many syntaxes from just one vendor called Daylight222www.daylight.com.
SMILES was introduced and subsequently standardised in OpenSMILES. CurlySMILES was introduced to annotate SMILES strings. CHUCKLES, its extension, CHORTLES, as well as BigSMILES were introduced to represent polymeric molecules [11]. SMIRKS was introduced as an extension of SMARTS to represent chemical reactions [31] – used in the Open Reaction Database. InChi V1, not part of Daylight, was introduced and can represent basic prototautemerism [32]. SMIRKS allows for string representation of reactions. It does not represent reaction features like temperature or pressure or catalysts. One feature it lacks is the ability to indicate the directionality of the reaction, which took a new language to be implemented ten years after the release of SMIRKS, called RinChi [11], [12]. This was further extended years later, with another language called ProcAuxInfo, which enables the storage of metadata relating to yields, temperature, concentration, and other reaction conditions. Although, this represents definite progress, the profusion of new syntaxes and grammars, and their creative backronyms, create avoidable complexity for the cheminformatician.
[33] introduces Balsa, a fully-specified language designed as a subset of SMILES. It aims to address the lack of a comprehensive SMILES specification and the resulting practical problems, such as the diversity of implementations, difficulties in standardization efforts, and challenges in developing extensions. Balsa provides a formal, machine-readable grammar that defines its syntax and describes semantics at three levels: constitution, configuration, and conformation. However, while Balsa improves specificity and compatibility, its compact representation may sacrifice some readability compared to more verbose alternatives. Additionally, like SMILES, Balsa strings can become complex and difficult to interpret for larger, more intricate molecules.
3.13 Fingerprints
@@
Fingerprints involve encoding key features of a molecule, typically into a vector representation and can offer convenience for machine learning applications. However, a significant drawback of fingerprints is their inability to consistently represent tautomers, which are molecules that have the same atomic composition but different structural arrangements. Tautomers often have distinct molecular structures, leading to different encodings in the bitmaps or fingerprints used for molecular database searching [34]. This inconsistency can result in the failure to identify the presence of a particular molecule in a database accurately.
@@
Moreover, while fingerprints can capture the presence of specific substructures within molecules, they are susceptible to bit collisions and vector sparsity, which limit their overall representational power [22]. Bit collisions occur when multiple distinct substructures are mapped to the same bit in the fingerprint, leading to ambiguity and potential false positives. Vector sparsity arises when the fingerprint representation is high-dimensional, but only a small fraction of the bits are actively used, resulting in inefficient memory usage and computational overhead.
@@
A fundamental issue with numeric vector representations, such as fingerprints, is the lack of inherent semantics associated with the individual numbers or floating-point values. Without additional contextual information or a predefined mapping, the numeric values in the vector do not carry intrinsic meaning. This lack of semantics may hinder the interpretability and reasoning capabilities of machine learning models trained on such representations.
@@
Furthermore, the type information associated with ADTs can guide and constrain the machine learning and software development process [^3] See Conor Mcbride talk about this issue: https://www.youtube.com/watch?v=v2yXrOkzt5w. The type constraints restrict the search space and focus learning algorithms on meaningful and plausible solutions. This may well lead to more efficient and effective learning, as the model can leverage the structure and semantics embedded in the ADT representation.
@@
Fingerprints are of course useful, but can be defined on the more general type data type by the programmer. For instance, given MAP4 touted as “One Molecular Fingerprint to rule them all” [35], can be constructed from an ADT, but fingerprints can lose information, cause bit collisions and lose algebraic properties of the elements of the fingerprint. These algebraic properties such as group theoretic properties can be used to predict molecular properties. Variational Auto Encoders (VAEs) also have problems, [6], whom claim that obtaining structured output to be a hard and currently unsolved problem, even for a simple graph data structure.
3.14 RDKit as representation
One might also ask whether existing and widely used Cheminformatics libraries support representation. Behemoth Cheminformatics programming libraries, such as RDKit, [10] and lightweight molecular manipulation libraries that exist on top of these e.g. datamol [^4] https://datamol.io/ and https://www.rdkit.org/ (which builds on NumPY [36], Pandas, RDKit and Python 3+), Datamol explicitly [^5] https://doc.datamol.io/stable/tutorials/Preprocessing.html don’t advocate themselves as providing a prototypical representation. These libraries parse representations to computational representations and then apply functions to them, before outputting them to strings like SMILES.
In the RDKit’s Mol classes, molecules are stored as a graph in the form of an adjacency list, with optional extra information in the form of properties and ring information. The internal representation therefore inherits the limitations of 2D graph based representations already discussed for computational purposes. Bonds in the graph can be labeled with 22 pre-defined bond types, and limited stereochemical features can be described, again with labeling.
The computational representation of RDKit in C++ is complex, taking up over 400 lines of highly complex code and the molecule class taking up 873 lines of code as of 2022. There are no clear invariants, or guarantees about the management of side effects in snippets shown below. The ubiquitous presence of null pointers and highly abstruse C++ keywords, scattered throughout the code, insinuates points of failure of the program. It is unclear, whether and when the program may crash and what the consequences might be. The following snapshot of lines $792$ to $821$ in ROMol.h [^6] https://www.rdkit.org/docs/cppapi/ROMol_8h_source.html which represents the molecule class, provides a brief glimpse into the extreme complexity of code:
1 #ifdef RDK_USE_BOOST_SERIALIZATION
[…]
3 //! @{
4 template <class Archive>
5 void save(Archive &ar, const unsigned int version) const;
6 template <class Archive>
7 void load(Archive &ar, const unsigned int version);
8 BOOST_SERIALIZATION_SPLIT_MEMBER()
9 //! @}
10 #endif
11
12 private:
13 MolGraph d_graph;
14 ATOM_BOOKMARK_MAP d_atomBookmarks;
15 BOND_BOOKMARK_MAP d_bondBookmarks;
16 RingInfo *dp_ringInfo = nullptr;
17 CONF_SPTR_LIST d_confs;
18 std::vector<SubstanceGroup> d_sgroups;
19 std::vector<StereoGroup> d_stereo_groups;
20 std::unique_ptr<boost::dynamic_bitset<>> dp_delAtoms = nullptr;
21 std::unique_ptr<boost::dynamic_bitset<>> dp_delBonds = nullptr;
22
23 friend RDKIT_GRAPHMOL_EXPORT std::vector<SubstanceGroup> &getSubstanceGroups(
24 ROMol &);
25 friend RDKIT_GRAPHMOL_EXPORT const std::vector<SubstanceGroup>
26 &getSubstanceGroups(const ROMol &);
27 void clearSubstanceGroups() { d_sgroups.clear(); }
28
29 protected:
30 unsigned int numBonds{0};
Listing 1: This is just a small fragment of the molecule class in RDKit. It demonstrates the sheer complexity and obfuscation of a concept that can be represented more simply in the ADT proposed in this article.
We aim to demonstrate that our ADT-based representation of molecules in Haskell offers a more concise, expressive, and type-safe approach. By leveraging the power of algebraic data types and the strong static typing of Haskell, we can define molecular structures in a declarative and compositional manner. The resulting code is more readable, maintainable, and less prone to runtime errors. The use of pure functions and the explicit management of side effects in a pure functional language such as Haskell, will enable easier reasoning about the behavior of the program and facilitate formal verification techniques.
Moreover, we claim that the ADT-based representation provides a clear separation of concerns between the data structure and the operations performed on it, interface versus implementation. This allows for a separation of concerns between those who are designing better algorithms for the data type and those who only need to use the data type. The use of higher-order functions and type classes in Haskell further enhances the reusability and composability of the code.
While cheminformatics libraries like RDKit and datamol have been widely adopted and have demonstrated their utility in various applications, their complex often OOP-based representations can hinder the development and maintenance of robust and reliable software. The proposed ADT-based approach offers a compelling alternative that prioritizes simplicity, type safety, and functional mathematical rigor.